OpenStudy (anonymous):

what iis the 210th term for the arithmetic sequence 3,8,13,18

OpenStudy (dls):

\[\LARGE a_{210}=a_1+(2n-1)d\] where n is the number of terms \[\LARGE a_{210}=3+(2(210)-1)5\] Do u get it? Can you solve further?

OpenStudy (dls):

d is the common difference between any two terms say 3.8 so it'll be 8-3=5 or take 13 and 8 13-8=5..etc